OpenStudy (anonymous):

Find the directional derivative of f(x,y)= sqrt(xy) at P(2,8) in the direction of Q(5,4)

OpenStudy (amistre64):

im gonna have to play with this just cause im outta practice :)

OpenStudy (anonymous):

ahahahhok, i love you amistre btw :) you save my life constantly

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcIII/DirectionalDeriv.aspx im reading up on it from this site :)

OpenStudy (anonymous):

yah i use that frequently

OpenStudy (amistre64):

so, vector from P to Q = Q-P 5,4 2,8 ---- <3,-4> since this is a 345 triangle I know the mag = 5 so the unit vector is: <3/5, -4/5> for that part

OpenStudy (amistre64):

I spose del F is next

OpenStudy (anonymous):

ok thats where i had trouble because i was not given a specific vector direction. So i need to make one from P to Q?

OpenStudy (amistre64):

\[\nabla f(x,y)= sqrt(xy)\] \[Fx = \frac{y}{2\sqrt{xy}}\] \[Fy= \frac{x}{2\sqrt{xy}}\]

OpenStudy (amistre64):

yes, from P in the direction of Q

OpenStudy (anonymous):

ok i think i can do the rest that was my main question. thanks again pal. I am just studying for a midterm i have in 3 hrs :)

OpenStudy (amistre64):

Do we then dot these together? <Fx,Fy> . <3/5,-4/5> ?

OpenStudy (amistre64):

midterms are fun lol

OpenStudy (anonymous):

they suck my left nut lol. ill be posting a few more questions soon if you could answer id appreciatge it!

OpenStudy (amistre64):

if im able :) ill try